Sixteen oil tins are taken at random from an automatic filling machine. The mean weight of the tins is 14.2 kg, with a standard deviation of 0.40 kg. Can we conclude that the filling machine is wasting oil by filling more than the intended weight of 14 kg, at a significance level of 5%? (5 marks)


1
Expert's answer
2022-06-16T09:50:55-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 14"

"H_1:\mu>14"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=15" and the critical value for a right-tailed test is "t_c =1.75305."

The rejection region for this right-tailed test is "R = \{t:t>1.75305\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{14.2-14}{0.40/\sqrt{16}}=2.5"


Since it is observed that "t=2.5>1.75305=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=15" degrees of freedom, "t=2.5" is "p=0.012253," and since "p=0.012253<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 14, at the "\alpha = 0.05" significance level.